OpenStudy (anonymous):

In terms of n, find the sum of the first n terms of the arithmetic sequence: -4, 5, 14, 23, 32, ..... can you show the steps??

Parth (parthkohli):

The standard way of writing a sum is \(\Large {n \over 2}(a_1 + a_n)\)

Parth (parthkohli):

So just substitute a1 with 14, and you're done :)

OpenStudy (anonymous):

so n/2 (-4 + 14)

OpenStudy (anonymous):

can you explain it to me?

OpenStudy (ash2326):

Do you know arithmetic sequence?

OpenStudy (anonymous):

ehh kind of, i just want to know how to write it out and solve it

OpenStudy (ash2326):

Do you know the relation to find the nth term of an arithmetic sequence?

OpenStudy (anonymous):

ahhhh nooo.

OpenStudy (ash2326):

Let the nth term be an, d the common difference and a1 be the first term. \(a_n\) is given as \[a_n=a_1+(n-1)d\] Now what's a1 and d here?

OpenStudy (anonymous):

a1 is -4 and would be +9

OpenStudy (anonymous):

right?

OpenStudy (ash2326):

Yeah, you're right:D Now sum of n terms of a sequence is given by the following \[S_n=n \times\frac{ (a_1+a_n)}{2}\]

OpenStudy (ash2326):

Can you plugin the values and write the relation ?

OpenStudy (anonymous):

what would \[a_{n}\] be?

OpenStudy (ash2326):

Use the relation that I posted earlier! \[a_n=a_1+(n-1)d\]

OpenStudy (anonymous):

\[a_{n} = -4 + 9n - 9 ?\]

OpenStudy (ash2326):

Yeah, you're right Now plugin this in Sum formula to find sum in terms of n

OpenStudy (anonymous):

\[S_{n} = n \times \left(\begin{matrix}-4 -4+9n-9 \\ 2\end{matrix}\right) ?\]

OpenStudy (ash2326):

\[S_n=n \times \frac{-4-4+9n-9}{2}\] Now simplify this:)

OpenStudy (ash2326):

Did you understand?

OpenStudy (anonymous):

no because i dont know what N is supposed to be

OpenStudy (ash2326):

We don't know n here, we have to leave the answer in terms of n. n is the no. of terms for which you want the sum

OpenStudy (anonymous):

i still don't understand

OpenStudy (ash2326):

Ok, :) Tell me which part you don't get